Brief Exercise 4-5 $84,600, respectively. The cost driver for each activity and the estimated annual usage are number of setups 2,800, machine hours 24,300, and number of inspections 1,800. Compute the overhead rate for each activity. Machine setupss Machining N8.sns Company identifestre ecthitr s oh asc manufe uring erocess: mactine sectuos, muchbnino, and inspections Ei bd anal veaheand olst for sc etivty 103016005413,100, and per setup per machine hour per inspection inspections Click if you would like to Show Work for this question: Open Show Work
Answer #1
answer
overhead rates for each activity
machine setups 72 per setup
machining 17 per machine hour
inspections 47 per inspection
computation of over head rate for each activity
overhead rate = estimated overhead cost/estimated level of cost driver activity
machine setups machining inspections
=201600/2800 =413100/24300 =84600/1800
72 17 47